#13 i have tired ti figure this out but I do not know where to start
The questions asked
A discount store sold plastic cups for 3.50 and each ceramic cups for 4 dollars each. If 400 cups were sold for a total of $1,458, how many cups of each type were sold ? what is the dollar value of each type of cup sold







Answer by stanbon(75887)

A discount store sold plastic cups for 3.50 and each ceramic cups for 4 dollars each. If 400 cups were sold for a total of $1,458, how many cups of each type were sold ? what is the dollar value of each type of cup sold
-------------------------------------------
Equations:
Quantity: p + c = 400
Value: 3.5p +4c = 1458
----------------------------
Multiply thru the 1st equation by 4 to set up the system for elimination:
4p + 4c = 1600
3.5p+4c = 1458
--------------------
Subtract 2nd Eq. from 1st to get:
0.5p= 142
p = 284 (# of plastic cups sold)
dollar value = 3.5*284 = $994.00
--------------------
Substitute into p+c=400 to solve for "c":
284 + c = 400
c = 116 (# of ceramic cups sold)
dollar value = $464
=====================
Cheers,
Stan H.